at soem point during my life i will be exactly 3 times older than my granddaughter please write the equasion and how I would do it. I am 57 and my grand is 4 Thank you so much!
Answer by sachi(548)

say after x yrs you will be exactly 3 times older than your granddaughter
so after x yrs you will be 57+x yrs & your grand will be 4+x yrs
so the eqn is 57+x=3(4+x)
ans


the soln will be
57+x=3(4+x)=12+3x
or 3x-x=57-12=45
or 2x=45
or x=22.5 yrs